10 mol of N2 gas expands at constant temperature of 300 K from 2 L at 8 atm to 6 L at 4 atm. How much work is performed by the gas?


1
Expert's answer
2021-02-17T10:40:13-0500

Since temperature is constant, process is isothermal. Work done in isothermal process is given by, "W = 2.303nRTlog(V_2\/V_1)"

Putting value, "W=2.303\times 10\times 8.3\times 300log(6/2)=27.36KJ"